OpenStudy (anonymous):

If i is raised to an even power, then it can not simplify to be Answers: -1, 1, i

OpenStudy (unklerhaukus):

\[i^0=1\\i^1=i\\i^2=-1\\i^3=-i\\\quad~\vdots\]

OpenStudy (anonymous):

so the answer is -1

OpenStudy (anonymous):

@bmx4life no, do out more iterations if you need. i^0=1 i^1=i i^2=-1 i^3=-i i^4=1 i^5=i i^6=-1 i^7=-i i^8=1 do you see the answer now?

OpenStudy (anonymous):

i

OpenStudy (anonymous):

correct! Good job :)

OpenStudy (anonymous):

ok thanks

OpenStudy (unklerhaukus):

if \(i\) is raised to a multiple fo two it will equal \(-1\) or \(1\) if \(i\) is raised to an multiple of four it will equal \(1\)